Fw: Shift report 20000619 16:00-24:00

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Tue Jun 20 2000 - 12:12:09 EDT

  • Next message: Flemming Videbaek: "Fw: Shift report 20000621 16:00-24:00"

    ----- Original Message ----- 
    From: http account <http@pii3.brahms.bnl.gov>
    To: <brahms-l@bnl.gov>; <olchansk@bnl.gov>; <videbaek@bnl.gov>
    Sent: Monday, June 19, 2000 11:00 PM
    Subject: Shift report 20000619 16:00-24:00
    
    
    > 
    > ------------------------------------------------------------
    > Date: Mon Jun 19 23:49:54 2000, Shift: 16:00-24:00
    > Supervisor: DB, with: YB
    > Log book#2 pages 41 - 45
    > 
    > Shift Summary:
    > ------------------------------------------------------------
    > 1. Beams cogged at gamma=70 for smoe time. 
    > Once for about 1.5 hours.
    > 2. Three runs 1440, 1441 and 1444. Several possible events 
    > candidates in Run1444
    > 3. T1 tripped once. Had to turn off MTPC1 because of the high
    > anode current (.8uA), most probably because of the beam 
    > scraping. RHIC's Blue/Yellow Lifitime monitor died several 
    > times, then came alive again.
    > 
    > ------------------------------------------------------------
    > 
    



    This archive was generated by hypermail 2b29 : Tue Jun 20 2000 - 11:06:36 EDT